need all of number 2 and three answered
Calculate the missing value in each problem. Delt
Answer #1

2

a)

dG =? dH = -217 dS = 151 T = 451

Since

G = H - TS

then

dG= -217 kJ/mol - 451K*(0.151 kJ/mol) = 285.1 kJ/mol

b)

dG =147 dH = dS = -98 T = 298

dG = dH - TdS

dH = dG + TdS = 147 kJ/mol + 298*-0.098kJ/mol = 117.8 kJ/mol

c)

dG =-20.4 dH =107 dS = 211 T = ?

dG = dH - TdS

T = -(dG-dH)/dS = -(-20.4 - 107)/0.211 = 603.8K

3)

m = 37.3 g of H2O

Cp = 4.184 J/gC

Tf = 49.9

Ti = 15.2

Q = ?

Q = m*Cp(Tf-Ti)= 37.3g*4.184J/gC*(49.9-15.2) = 5415.4 J